Determine the molecular formula of butane given the molecular mass of 58.12 g/mol and empirical formula of C2H5.
1
Expert's answer
2021-10-07T03:48:03-0400

"Molecular formula= n (E.F)"

C2H5 =29

n=58/29 =2

M.F= C4H10